use method of substitution to solve the system of linear equations
1. 4a= 6 - b
+6 = +6
4a+6=b (stuck from here)

2 4a= 5 - b

Answer by Alan3354(69443)

use method of substitution to solve the system of linear equations
1. 4a= 6 - b
+6 = +6
4a+6=b (stuck from here)
2 4a= 5 - b
=============
If your 2 eqns are
use method of substitution to solve the system of linear equations
4a = 6 - b
4a = 5 - b
------------------
Sub for 4a
6 - b = 5 - b
6 = 5
Inconsistent, no solution.